OpenStudy (candycane):

can someone help me solve this using elimination? 5x-y+z=4 x+2y-z=5 2x+3y-3z=5

OpenStudy (candycane):

@Destiny_forever

OpenStudy (candycane):

@mathmate

OpenStudy (candycane):

I've watched all the stuff my teacher has asked us to and I still dont understand

OpenStudy (anonymous):

hint: "add" the first two equations together. that will eliminate the "z" variable

OpenStudy (candycane):

I did that my final thng that i got was 4x-y= 9 but after that ig et stuck

OpenStudy (aihberkhan):

I can help! :)

OpenStudy (aihberkhan):

Okay. Lets first use the first equation. Lets add the y and subtract 5x from both sides. When we do this we get: \[z = 4 + y - 5x\]

OpenStudy (candycane):

wait wouldnt it become 4 x?

OpenStudy (aihberkhan):

Yes! Good! I was just going step by step. But good job!

OpenStudy (candycane):

Also did you bring the z to the oppiste side of the equal side? if so why???

OpenStudy (aihberkhan):

Well, It depends on how you solve it. But you always want the variable you want to solve for first on a seperate side than anything else.

OpenStudy (aihberkhan):

So when you solve the three equations in such a way you should get: \[x = 1\] \[y = 3\] \[z = 2\]

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @Candycane

OpenStudy (candycane):

okay. But how do I solve to get those numbers??

OpenStudy (candycane):

This is for extra credit to bring up my grade and it is required to show all the steps, I'd ask my teacher but i wont hear from her for two weeks due to winter break

OpenStudy (candycane):

Nope.. Im still slightly stuck...

OpenStudy (candycane):

@Zarkon

OpenStudy (candycane):

@Agl202

OpenStudy (xmissalycatx):

I only have a few minutes so I'll do my best to help you... You have two equations.. \[5x-y+z=4\] \[x+2y-z=5\] So just find something that is easy to eliminate first... like z because they are the same value, just opposite signs (which is what you want so they can be crossed out). Now you're left with 5x - y = 4 and x + 2x = 5

OpenStudy (xmissalycatx):

Now you're going to do the exact same thing with the 2nd and last equations (because the first time we did it, it was the first and second equations..) So x + 2y - z = 5 and 2x + 3y - 3z = 5 you HAVE to eliminate z again since you did it the first time.. so in order to make it where you can eliminate it, you need to multiply the whole equation by -3.. so -3x - 6y + 3z = -15 and 2x + 3y - 3z = 5 which gives you -x - 3y = -10 and the very first equation you did was 4x + y = -1

OpenStudy (xmissalycatx):

Let me correct myself.. the very first one would end up to be 6x + y = 9 lol

OpenStudy (xmissalycatx):

continue the process I have to go.